Show that x = y = z+1 is a secant line of the sphere x^2 +y^2 +z^2 −x−y+z−1 = 0.
Also find the intercept made by the sphere on the line.
1
Expert's answer
2020-02-20T08:30:46-0500

Let assume "x=y=z+1=k"

now putting the value of x, y and z in the above equation of the sphere we get

"k^2+k^2+(k-1)^2-k-k+k-1-1=0
\newline
3k^2-3k-1=0"

k"=1.26 \ \& -0.26"

so we get two points where the line cuts the sphere :

"(1.26,1.26,0.26)\ and \ (-0.26,-0.26,-1.26)"

distance between these points is "\sqrt{(1.26+0.26)^2(1.26+0.26)^2(-1.26-0.26)^2}=2.63"


radius of the above sphere = "\sqrt{(0.5)^2+(0.5)^2+(0.5)^2+1}=1.32"

diameter of the sphere = "s\times radius =2\times1.32=2.64"


distance between the two points is less than the diameter of the sphere hence the given line is a secant the sphere and the intercept made by the line to the sphere is 2.63